OpenStudy (anonymous):

Joshua is creating a series of rectangles, one inside the other. The length of the first rectangle, Rectangle A, is five inches. Joshua dilates Rectangle A about its center by a scale factor of two to get Rectangle B. He repeats this process using the same scale factor till he gets a rectangle of length 80 inches. How many dilations did Joshua perform including the dilation of Rectangle A? 4 2 8 16

OpenStudy (anonymous):

i think the answer is 4, i am not sure though

OpenStudy (anonymous):

i agree... 4... first dilation: 5*2=10 second: 10*2 = 20 third: 20*2 = 40 fourth: 40*2 = 80

OpenStudy (anonymous):

ok thank you :)